Sketch the graph of the parabola.
y=-x^2+4x-7
I found the vertex which is (2,-3) so I graphed that point. What is the other point that I need to graph? I don't know how to get the other point in order to complete the graph of the parabola. Can you please include the other point and how you got the other point, thanks.

Answer by jim_thompson5910(35256)

Plug in any x value you want. I'm going to plug in something close to 2 (since the x coordinate of the vertex is x = 2)

So I'll pick x = 3

Plug in x = 3

y=-x^2+4x-7

y=-(3)^2+4(3)-7

y = -9 + 4(3) - 7

y = -9 + 12 - 7

y = -4

So when x is x = 3, y is y = -4.

So another point on the parabola is (3,-4)

By symmetry (along the axis of symmetry x = 2), another point is (1,-4). This is because x = 1 is exactly the same distance away from the axis of symmetry as x = 3 is. So these two x values produce the same y value.

You can plug in any amount of x values you want to generate any number of extra pairs of points.